OpenStudy (anonymous):

I have a bunch of problems like this and I need help understanding how to do them (5)^-5(5)^7 My choices are A. 1/25 B. 25 C. 12 D. 2

OpenStudy (whpalmer4):

\[5^{-5}*5^7\]

OpenStudy (whpalmer4):

First, do you understand what the notation means?

OpenStudy (anonymous):

we'll for the sake of my confusion go with no just to get an explanation

OpenStudy (whpalmer4):

Okay. Think of the exponent (the little number up top) as a shorthand. \[x^2 = x*x\]\[x^3=x*x*x\]\[x^4=x*x*x*x\]\[x^{537962} = x*x*x*x*\text{just kidding!}\] Okay so far?

OpenStudy (anonymous):

so far so good

OpenStudy (whpalmer4):

Now, if we have a negative exponent, that's the same as: \[x^{-2} = \frac{1}{x^2} = \frac{1}{x*x}\]\[x^{-3} = \frac{1}{x^3} = \frac{1}{x*x*x}\]and in general, \[x^{-n} = \frac{1}{x^n}\]and\[\frac{1}{x^{-n}} = x^{n}\]

OpenStudy (anonymous):

ok I've got you so far.

OpenStudy (whpalmer4):

Now, the property of exponents this problem is apparently to exercise is this one: \[x^{n}*x^{m} = x^{n+m}\] In other words, when you multiply exponents \(\bf\text{and they have the same base}\) you add the exponents to get the exponent of the product. Examples: \[2^2*2^3 = (2*2)*(2*2*2) = 2*2*2*2*2 = 2^{2+3} = 2^5\]Checking:\[2^2=4, 2^3 = 8, 4*8 = 32, 2^5 = 2*2*2*2*2= 32\] \[2^{-2}*2^4 = \frac{1}{2^2}*2^4 = \frac{1}{2*2}*(2*2*2*2) = 2*2 = 4\]\[2^{-2}*2^2 = 2^{-2+4} = 2^2 = 4\]

OpenStudy (whpalmer4):

sorry, I mistyped the last line, should be \[2^{-2}*2^4 = 2^{-2+4} = 2^{2} = 4\]

OpenStudy (whpalmer4):

two other things to remember are that \[x^1 = x\]and \[x^0=1\](the latter being true so long as \(x\ne0\), if \(x = 0\) it is undefined)

OpenStudy (whpalmer4):

So your problem is \[5^{-5}*5^{7} = \]what is the answer in exponential form, after you've added the exponents?

OpenStudy (anonymous):

5^2?

OpenStudy (whpalmer4):

Yes! Now, if you convert that to a number such as found in your answer choices, which one would it be?

OpenStudy (anonymous):

A.?

OpenStudy (whpalmer4):

What does \(5^2\) mean?

OpenStudy (anonymous):

mistype sorry I meant B.

OpenStudy (whpalmer4):

That's correct. \[5^2 = 5*5 = 25\]

OpenStudy (whpalmer4):

Here, I'll give you another problem with an answer from that list: \[5^1*5^{-3} = \]

OpenStudy (anonymous):

Thanks for the Help I understand what im doing now :)

OpenStudy (anonymous):

erm lets 5^-2= 25?

OpenStudy (whpalmer4):

Well, the first part of the answer is correct. \[5^1*5^{-3} = 5^{1-3} = 5^{-2}\]Remember, if you have a negative exponent, put it on the other side of the division bar and make it a positive exponent: \[5^{-2} = \frac{1}{5^2} = \]

OpenStudy (anonymous):

oh so 1/25

OpenStudy (whpalmer4):

Right! There's only 1 number for which \[x^n = x^{-n}\]Can you guess what it might be?

OpenStudy (anonymous):

2?

OpenStudy (whpalmer4):

Hint: it appears in the question :-)

OpenStudy (anonymous):

I give whats the answer lol?

OpenStudy (whpalmer4):

It's not 2: say \(n = 3\), then \[2^3 = 8\]but \[2^{-3} = \frac{1}{2^3} = \frac{1}{8}\]and that isn't the same as \(8\), I hope you'll agree! How about 1? \[1^n = 1*1*1*1*...*1=1\] \[1^{-n} = \frac{1}{1^n} = \frac{1}{1*1*1*1*...*1} = \frac{1}{1} = 1\]

OpenStudy (anonymous):

so 1 I take it lol

OpenStudy (whpalmer4):

Yes. 1 is the only number that works. A gentle way of suggesting that if the answer to one problem is \(x^n =\text{some number}\) then you better not get \(x^{-n} = \text{same number}\) unless \(x = 1\)...

OpenStudy (anonymous):

Well thank you whpalmer I have repaid you with the medals and a fan :)

OpenStudy (whpalmer4):

Hopefully that is enough to get you through the current homework with flying colors :-)